OpenStudy (anonymous):

A = 41°, B =61°, c = 19 solve using the law of sines? show steps please so I can actually learn how to do this! Thank you!:) also, please round to the nearest tenth if possible. Thanks again. :)

OpenStudy (johnweldon1993):

Alright so the law of sines \[\large \frac{a}{\sin A}= \frac{b}{\sin B} =\frac{c}{\sin C}\] where a , b and c are the sides of your triangle and A, B and C are the angles of your triangle What would we do first? do you know?

OpenStudy (anonymous):

Nope. honestly i have no idea 

OpenStudy (johnweldon1993):

Dont worry about it, we'll go through it step by step...

OpenStudy (anonymous):

wait, like 41/sinA 61/sinB and 19/sinC?

OpenStudy (johnweldon1993):

So what you have KNOWN so far is A = 41 degrees B = 61 degrees and c = 19 if we substituted all this into that formula up there we would have \[\large \frac{a}{\sin 41}= \frac{b}{\sin 61} =\frac{19}{\sin C}\] What we want to do...is complete 1 of these ratios that we can compare the others to...

OpenStudy (anonymous):

oh, okay, thanks. so how do i do that? lol

OpenStudy (johnweldon1993):

you were close...just remember the A, B and C are the angles...and go on the bottom...and the a, b and c are the side lengths and go on top...

OpenStudy (johnweldon1993):

So now...you have 2 of the 3 angles in a triangle right? What do the angles of a triangle sum up to?

OpenStudy (anonymous):

180, so 41+61= 102. so angle 3 is 78?

OpenStudy (johnweldon1993):

That is correct! so now we have \[\large \frac{a}{\sin 41}= \frac{b}{\sin 61} =\frac{19}{\sin 78}\]

OpenStudy (johnweldon1993):

Now that you have 1 completed ratio...you can use it to find the missing variables...for example...

OpenStudy (johnweldon1993):

\[\large \frac{b}{\sin 61} =\frac{19}{\sin 78}\] and solve for b...

OpenStudy (anonymous):

and how do I do that?

OpenStudy (johnweldon1993):

Well you want to isolate b right? so multiply both sides of this equation by sin61 \[\large b = \frac{19\times \sin 61}{\sin 78}\] throw this in your calculator and solve for b...what do you get?

OpenStudy (anonymous):

uhmm i got -12.etc etc so i think i did it wrong. i couldnt get the first part of it to go over the second. if that makes sense?

OpenStudy (johnweldon1993):

make sure you calculator is in "degree" mode

OpenStudy (anonymous):

uhm my calculator is on my phone and doesnt have options like that.:/

OpenStudy (johnweldon1993):

Even if you turn it to the side? well regardless you get... https://www.google.com/#q=(19*sin(61degrees))%2F(sin(78degrees))

OpenStudy (johnweldon1993):

And remember to round to the nearest tenth so that comes to b = 17

OpenStudy (anonymous):

lol no it turns scientific if i do that though. & oh okay thanks. so now what?

OpenStudy (johnweldon1993):

So now you have to solve for a...you have 'b' and sinB....so use the ratio you just solved and use it to solve for 'a' the same way we just did for 'b'

OpenStudy (johnweldon1993):

\[\large \frac{a}{\sin 41} =\frac{17}{\sin 61}\] and solve for 'a'

OpenStudy (anonymous):

so a/sin41*19/sin78?

OpenStudy (johnweldon1993):

Yes you could also do that @PitbullsAreForever ...since every ratio is equal... but remember it is = not *

OpenStudy (anonymous):

oh so then i will multiply 78 by 19? or??im lost i think :/

OpenStudy (wolf1728):

It can all be done without using the law of sines: sin (41°) = hgt / 19 hgt = sin (41°) * 19 hgt = 12.465 side AD² = 19² - 12.465² side AD = 14.34 tan 78° = 12.465 / side DC side DC = 12.465 / 4.70 side DC = 2.65 side b = 14.34 + 2.65 = 16.99 side a = 12.465² + 2.65² side a = 12.74